Grade 11Physical Chemistry

A compound has N=25.94% &O = 74.06%. If it`s vapour density is 54.2, it`s molecular formula is what?

Krissh
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Dear Student,
For assurance of sub-atomic recipe most importantly number of gram molecules are discover 
Number of gram particles of Nitrogen= 25.94/14 = 1.85 
Number of gram particles of Oxygen= 74.06/16 = 4.62 
Nuclear proportion between N:O by partitioning the littlest proportion 
1.85/1.85:4.62/1.85= 1:2.49 
Change over the part number into entire number appropriate digit is duplicated with nuclear proportion that is 2 
N:O=2(1:2.49)= 2:5 
In this manner the atomic equation is N2O5(Dinitrogen pentaoxide).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)